OpenStudy (anonymous):

FAN & MEDAL to who helps me solve: 2f(5f-2)-10(f^2-3f+6)=-8f(f+4)+4(2f^2-7f)

OpenStudy (anonymous):

2f(5f−2)−10(f^2−3f+6)=(−8f)(f+4)+4(2f^2−7f) (2f)(5f)+(2f)(−2)+(−10)(f2)+(−10)(−3f)+(−10)(6)=(−8f)(f)+(−8f)(4)+(4)(2f2)+(4)(−7f)(Distribute) 10f^2+−4f+−10f^2+30f+−60=−8f^2+−32f+8f^2+−28f

OpenStudy (anonymous):

(10f^2+−10f^2)+(−4f+30f)+(−60)=(−8f^2+8f^2)+(−32f+−28f)(Combine Like Terms)

OpenStudy (anonymous):

\[\huge~\rm~26f−60=−60f \] \[\huge~\rm~26f−60+60f=−60f+60f \] Add 60 to both sides. \[\huge~\rm~86f−60+60=0+60\] \[\huge~\rm~86f=60\]

OpenStudy (anonymous):

Divide both sides by 86. \[\huge~\rm~86f/86=60/86\]

OpenStudy (anonymous):

\[\huge~~\rm~f=? \]

OpenStudy (anonymous):

Hope i didnt make is too confusing D:

OpenStudy (anonymous):

very confusing , sorry. but i combined -4f and 30f but i dont see it in your explanation. @preciouspearl